# CAMPBELL & REMIRO-AZÓCAR

### v1

@@ 4.3 Estimands
The estimand of interest is the ATC, as defined in Section 2. We adopt the logit link function $g(p)=\ln(p/(1-p))$ for marginal outcome probability $p$, such that the ATC is on the marginal log-odds ratio scale. The values of the ATC estimands were calculated numerically, by simulating 1 million binary outcomes using the true data-generating mechanisms outlined in Section 4.2. Data-generating mechanisms KS1, KS2, KS3 and KS4 correspond to true ATCs of 1.128, 1.214, 1.053 and 1.163, respectively.
@@ 4.4 Methods
  2. 2.
    The IOW estimator with weights derived using the “modeling” approach, as per Equation 3 (Section 3.2).
  3. 3.
    The IOW estimator with normalized weights derived using the “modeling” approach, as per Equation 4 (Section 3.2).
  4. 4.
@@
  5. 5.
    The G-computation estimator, as per Equation 19 (Section 3.4).
  6. 6.
    The doubly robust (DR) augmented weighting estimator with “modeling” IOW weights, as per Equation 21 (see Section 3.4).
  7. 7.
    The doubly robust (DR) augmented weighting estimator with normalized “modeling” IOW weights, as per Equation 23 (see Section 3.4).
  8. 8.
    The doubly robust (DR) augmented weighting estimator with MAIC (entropy balancing) weights, as per Equation 25 (our main contribution, see Section 3.4).
  9. 9.
    The augmented “weighted G-computation” estimator described in Section 3.5, said to be doubly robust for the ATE where the outcome model is a GLM with a canonical link function, using the (normalized) “modeling” IOW weights.
  10. 10.
    The augmented “weighted G-computation” estimator described in Section 3.5, but using the MAIC (entropy balancing) weights instead of the “modeling” IOW weights.
For all estimators, we computed 95% confidence intervals using the non-parametric bootstrap approach described in Section 3.6. Specifically, we used $B=100$ bootstrap resamples of the concatenated SAT and external control IPD to approximate the standard error of $\widehat{\textrm{ATC}}$ and subsequently constructed Wald-type confidence intervals.
@@ 4.5 Performance measures
To assess the performance of the estimators in our simulation study, we computed several key metrics: bias, empirical standard error (ESE), and 95% confidence interval (CI) coverage. Bias was calculated as the difference between the average of the point estimates across simulations and the true estimand value, providing a measure of systematic error. The ESE was computed as the standard deviation of the point estimates across simulations, reflecting the precision or variability of the different estimators. The 95% CI coverage was determined as the proportion of simulated datasets in which the constructed 95% confidence interval contained the true estimand value, evaluating the quality of interval estimation. In addition, we estimated Monte Carlo standard errors (MCSEs) using the formulas provided by Morris et al (2019) to quantify the uncertainty in the performance measures due to using a finite number of simulations.91
@@ 4.6 Results
The complete results for the simulation study are displayed in Tables 1-4. Note that the MCSEs are less than 0.001 for bias, ESE and 95% CI coverage, and considered negligible for all three performance metrics. Consequently, they have been omitted from Table 1 to Table 4.
As expected, the naïve estimator is biased in all four simulation scenarios. Under $n=1000$, the G-computation estimator is virtually unbiased in Scenario KS1 (bias = 0.003; see Table 1) and Scenario KS3 (bias = 0.004; see Table 3), where the outcome model is correctly specified; but exhibits bias where the outcome model is misspecified, in Scenario KS2 (bias = 0.067; see Table 2) and, particularly, in Scenario KS4 (bias = 0.516; see Table 4). Under $n=1000$, the IOW estimators based on modeling weights are virtually unbiased in Scenario KS1 and Scenario KS2, where the propensity score model for data source assignment is correctly specified; but are biased in Scenario KS3 and Scenario KS4, where the propensity score model is misspecified. Notably, in KS3 where the propensity score model is misspecified, the IOW estimator with normalized weights displays a bias in the opposite direction (and of a lesser magnitude) than the version with non-normalized weights.
While, under $n=1000$, the non-augmented MAIC (entropy balancing) estimator is unbiased in Scenarios KS1 and KS2, it exhibits bias in Scenarios KS3 and KS4, which confirms its lack of double robustness where the outcome-generating model is logistic. This motivates the need for explicit augmentation with an outcome model. All the augmented estimators appear to be unbiased with $n=1000$ where only one of either the propensity score model or the outcome model is correctly specified (Scenarios KS2 and KS3), but not both (Scenario KS4). This suggests that these estimators do have the doubly robust property and increased bias-robustness compared to MAIC, providing two opportunities for valid inference and additional protection against model misspecification.
@@
The trends observed for the bias are similar under $n=200$ with the caveat that the (augmented and non-augmented) weighting estimators that appeared unbiased under $n=1000$, exhibit some small-sample bias in the corresponding scenarios. This is particularly notable in Scenario KS2 and, to a lesser extent, in Scenario KS1, and is probably due to small effective sample sizes after weighting.
@@
Under correct specification of the outcome model (KS1 and KS3), G-computation is the most precise covariate-adjusted estimator, but the augmented estimators are almost as precise (e.g., compare the G-computation estimator which obtains ESE = 0.150 to the DR with MAIC weights estimator which obtains ESE = 0.169 for KS1 with $n=1000$; see Table 1), and generally produce precision gains versus their respective non-augmented weighting counterparts. When both the outcome model and the propensity score model are correctly specified (KS1), all augmented estimators have increased precision compared to the non-augmented weighting estimators based on modeling weights, but not necessarily against MAIC (any increase in precision for $n=1000$ is modest).
When only the outcome model is correctly specified (KS3), all augmented estimators display greater precision than the non-augmented weighting estimators, including MAIC. When only the propensity score model is correctly specified (KS2), outcome model misspecification does not induce a loss in precision for the augmented estimators compared to their non-augmented weighting counterparts, with the exception of IOW-based estimators with normalized modeling weights. All augmented estimators have similar precision to G-computation in KS2; albeit with $n=200$, the precision gains for G-computation are more notable, probably due to limited effective sample sizes after weighting.
@@
As expected, when both the propensity score model and the outcome model are misspecified, all estimators are biased. Augmentation via an outcome model does not protect against the simultaneous misspecification of two models. There have been some concerns in the literature about doubly robust augmented estimators amplifying bias and variance when misspecified weights are combined with a misspecified outcome model.85 Such amplification does not take place in our simulation study; under dual model misspecification (KS4), the augmented estimators offer less bias and increased precision compared to their non-augmented weighting counterparts, and our proposed augmented MAIC estimator is the least biased of all estimators and the most precise of the augmented and non-augmented weighting estimators.
@@
$n=200$
1. The naïve estimator 0.618 0.328 0.539 1.292
2. IOW with weights from modeling 0.024 0.528 0.939 1.978
3. IOW with normalized weights from modeling 0.049 0.456 0.944 1.763
4. MAIC 0.033 0.420 0.959 2.241
5. G-computation 0.016 0.350 0.955 1.430
6. DR with “modeling” IOW weights 0.029 0.421 0.954 1.667
7. DR with normalized “modeling” IOW weights 0.029 0.414 0.948 1.610
8. DR with MAIC weights 0.029 0.412 0.953 1.713
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.027 0.404 0.940 1.583
10. Augmented “weighted G-computation” with MAIC weights 0.026 0.406 0.943 1.740
$n=1000$
1. The naïve estimator 0.604 0.143 0.009 0.561
2. IOW with weights from modeling 0.009 0.205 0.950 0.806
3. IOW with normalized weights from modeling 0.010 0.196 0.941 0.750
4. MAIC 0.006 0.171 0.942 0.659
5. G-computation 0.003 0.150 0.949 0.592
6. DR with “modeling” IOW weights 0.005 0.174 0.946 0.675
7. DR with normalized “modeling” IOW weights 0.005 0.174 0.946 0.666
8. DR with MAIC weights 0.005 0.169 0.941 0.651
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.004 0.169 0.942 0.648
10. Augmented “weighted G-computation” with MAIC weights 0.004 0.169 0.940 0.649
Table 1: Results from Scenario KS1, where both the outcome model and the propensity score model are correctly specified.
@@
$n=200$
1. The naïve estimator 0.223 0.322 0.910 1.275
2. IOW with weights from modeling 0.022 0.665 0.929 2.386
3. IOW with normalized weights from modeling 0.052 0.515 0.938 1.954
4. MAIC 0.052 0.501 0.960 2.747
5. G-computation 0.081 0.436 0.951 1.731
6. DR with “modeling” IOW weights 0.043 0.542 0.947 2.100
7. DR with normalized “modeling” IOW weights 0.043 0.520 0.941 2.003
8. DR with MAIC weights 0.039 0.490 0.950 2.044
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.049 0.480 0.936 1.875
10. Augmented “weighted G-computation” with MAIC weights 0.028 0.480 0.946 2.198
$n=1000$
1. The naïve estimator 0.220 0.141 0.665 0.556
2. IOW with weights from modeling 0.012 0.277 0.946 1.077
3. IOW with normalized weights from modeling 0.007 0.221 0.938 0.837
4. MAIC 0.006 0.205 0.934 0.777
5. G-computation 0.067 0.188 0.936 0.734
6. DR with “modeling” IOW weights 0.005 0.226 0.941 0.865
7. DR with normalized “modeling” IOW weights 0.006 0.224 0.937 0.848
8. DR with MAIC weights 0.005 0.205 0.936 0.775
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.009 0.202 0.935 0.778
10. Augmented “weighted G-computation” with MAIC weights 0.005 0.203 0.935 0.769
@@
$n=200$
1. The naïve estimator -0.033 0.301 0.952 1.208
2. IOW with weights from modeling 0.132 0.568 0.961 2.212
3. IOW with normalized weights from modeling -0.032 0.386 0.951 1.534
4. MAIC 0.121 0.346 0.955 1.518
5. G-computation 0.007 0.284 0.959 1.175
6. DR with “modeling” IOW weights 0.018 0.340 0.961 1.398
7. DR with normalized “modeling” IOW weights 0.017 0.328 0.957 1.335
8. DR with MAIC weights 0.015 0.310 0.956 1.290
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.012 0.302 0.954 1.242
10. Augmented “weighted G-computation” with MAIC weights 0.010 0.302 0.958 1.283
$n=1000$
1. The naïve estimator -0.040 0.134 0.937 0.528
2. IOW with weights from modeling 0.117 0.228 0.968 0.918
3. IOW with normalized weights from modeling -0.046 0.165 0.938 0.645
4. MAIC 0.104 0.146 0.889 0.573
5. G-computation 0.004 0.122 0.952 0.487
6. DR with “modeling” IOW weights 0.007 0.142 0.947 0.559
7. DR with normalized “modeling” IOW weights 0.007 0.140 0.946 0.549
8. DR with MAIC weights 0.006 0.132 0.946 0.517
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.005 0.127 0.949 0.505
10. Augmented “weighted G-computation” with MAIC weights 0.005 0.128 0.948 0.504
Table 3: Results from Scenario KS3, where the outcome model is correctly specified and the propensity score model is incorrectly specified.
In Section 3.5, we described an augmented “weighted G-computation” estimator that is doubly robust where the target estimand is the ATE and the outcome model is a GLM with a canonical link function. The results of the simulation study suggest that this estimator is also doubly robust for the ATC, noting that the employed logistic outcome model has a canonical link function (the logit link for a binomial outcome distribution). Interestingly, this augmented “weighted G-computation” estimator offers very similar performance, both in terms of bias and precision, than our proposed doubly robust augmented estimator with MAIC weights, regardless of whether (normalized) modeling IOW weights or MAIC weights are used. These three estimators are generally the least biased and most precise of all the augmented and non-augmented weighting estimators. Their similar performance is consistent with theoretical investigations by Sloczynski et al (2023),93 which suggest that the different versions of augmented weighting estimators are numerically equivalent when balancing weights are used and the target estimand is the ATT or the ATC.aaaIt is worth noting that the augmented weighting estimators investigated by Sloczynski et al (2023) use a different type of balancing weights, those from Imai and Ratkovic (2014),94 and a linear outcome model.93
@@
$n=200$
1. The naïve estimator 0.519 0.338 0.684 1.329
2. IOW with weights from modeling 0.800 0.783 0.908 2.763
3. IOW with normalized weights from modeling 0.573 0.475 0.757 1.832
4. MAIC 0.608 0.469 0.787 2.013
5. G-computation 0.532 0.383 0.745 1.544
6. DR with “modeling” IOW weights 0.576 0.459 0.767 1.831
7. DR with normalized “modeling” IOW weights 0.571 0.443 0.750 1.753
8. DR with MAIC weights 0.513 0.414 0.774 1.664
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.541 0.420 0.761 1.678
10. Augmented “weighted G-computation” with MAIC weights 0.540 0.429 0.781 1.780
$n=1000$
1. The naïve estimator 0.497 0.146 0.071 0.574
2. IOW with weights from modeling 0.788 0.359 0.334 1.425
3. IOW with normalized weights from modeling 0.520 0.199 0.249 0.765
4. MAIC 0.552 0.192 0.165 0.738
5. G-computation 0.516 0.162 0.104 0.635
6. DR with “modeling” IOW weights 0.542 0.188 0.178 0.729
7. DR with normalized “modeling” IOW weights 0.541 0.185 0.170 0.716
8. DR with MAIC weights 0.487 0.173 0.188 0.665
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.530 0.178 0.148 0.710
10. Augmented “weighted G-computation” with MAIC weights 0.539 0.183 0.153 0.708
@@
Assuming unbiasedness, interval estimation is appropriate if the coverage is approximately equal to 0.95; poor coverage can arise due to bias or to inadequate variance/interval estimation. Coverage is generally close to 0.95 for all covariate adjustment methods in the simulation scenarios under which they are unbiased, which suggests that our proposed non-parametric bootstrap approach for variance estimation is adequate. In the cases in which the covariate adjustment methods are unbiased, there is slight undercoverage for $n=1000$, with coverage rates between 0.934 and 0.950 across scenarios. Interestingly, coverage rates seem to increase for these methods under $n=200$ despite the small-sample bias, lying between 0.938 and 0.961 (except IOW with modeling weights in KS2 with 0.929), probably due to overly wide interval estimates due to small sample sizes. Note that, due to computational limitations, the non-parametric bootstrap approach in the simulation study was conducted with only $B=100$ resamples. This may have impacted the observed coverage rates and we suspect that coverage might be more appropriate when using a larger number of resamples.
While some covariate adjustment methods display bias-induced undercoverage in the scenarios under which they are biased (e.g., MAIC in KS3 under $n=1000$ or all estimators in KS4), they may also display adequate coverage (e.g., IOW with normalized modeling weights under $n=200$) because of excessively large standard errors, probably due to low effective sample sizes after weighting. As observed for KS4, bias-induced undercoverage tends to worsen with higher sample sizes, as interval estimates around the wrong target value become narrower. The naïve estimator displays discernible undercoverage in KS1, KS2 and KS4 (particularly under $n=1000$), not only due to bias but also due to overprecise standard errors that do not account for covariate differences.

### v2

@@ 4.3 Estimands
The estimand of interest is the ATC, as defined in Section 2. We adopt the logit link function $g(p)=\ln(p/(1-p))$ for marginal outcome probability $p$, such that the ATC is on the marginal log-odds ratio scale. The values of the ATC estimands were calculated numerically, by simulating 10 million binary outcomes using the true data-generating mechanisms outlined in Section 4.2. Data-generating mechanisms KS1, KS2, KS3 and KS4 correspond to true ATCs of 1.116, 1.215, 1.068 and 1.181, respectively.
@@ 4.4 Methods
  • 2.
    The IOW estimator as per Equation 3 (Section 3.2), with weights derived using the “modeling” approach with a logistic regression data source assignment model as per Equation 2.
  • 3.
    The normalized IOW estimator with normalized weights as per Equation 4 (Section 3.2) derived using the “modeling” approach with a logistic regression data source assignment model as per Equation 2.
  • 4.
@@
  • 5.
    The G-computation estimator, as per Equation 16 (Section 3.4) with the outcome model defined with a logistic link function.
  • 6.
    The DR augmented IOW estimator, as per Equation 18 (see Section 3.5) with the outcome model defined with a logistic link function.
  • 7.
    The DR augmented normalized IOW estimator, as per Equation 20 (see Section 3.5) with the outcome model defined with a logistic link function.
  • 8.
    The DR augmented MAIC (augmented entropy balancing) estimator, as per Equation 22 (our main contribution, see Section 3.5) with the outcome model defined with a logistic link function.
  • 9.
    The weighted G-computation (normalized IOW weights) estimator described in Section 3.6 with the outcome model defined with a logistic link function, and using the (normalized) “modeling” IOW weights.
  • 10.
    The weighted G-computation (MAIC weights) estimator described in Section 3.6 with the outcome model defined with a logistic link function, but using the MAIC (entropy balancing) weights instead of the “modeling” IOW weights.
  • 11-16.
    Estimators 11-16 are the same as Estimators 5-10 but using a non-canonical Cauchit link function, $q(x)=\textrm{tan}(\pi(x-0.5))$. See Morgan and Smith (1992)morgan1992note for an example of model fitting with the Cauchit.
To be clear, for Estimators 1-10, both the outcome model and the propensity score model for data source assignment will be correctly specified in KS1. For Estimators 11-16, the propensity score model for data source assignment will be correctly specified in KS1, but not the outcome model. In KS2, the outcome model will be incorrectly specified for all estimators fitting an outcome model (Estimators 5-16). Finally, in KS3 the propensity score model will be incorrectly specified for all estimators considering a propensity score model, and in KS4 both the outcome model and the propensity score model will be incorrectly specified.
For all estimators, we computed 95% CIs using the non-parametric bootstrap approach described in Section 3.7. Specifically, we used $B=200$ bootstrap resamples of the concatenated SAT and external control IPD to approximate the standard error of $\widehat{\textrm{ATC}}$ and subsequently constructed Wald-type CIs.
@@ 4.5 Performance measures
To assess the performance of the estimators in our simulation study, we computed several key metrics: bias, empirical standard error (ESE), and 95% CI coverage. Bias was calculated as the difference between the average of the point estimates across simulations and the true estimand value, providing a measure of systematic error. The ESE was computed as the standard deviation of the point estimates across simulations, reflecting the precision or variability of the different estimators. The 95% CI coverage was determined as the proportion of simulated datasets in which the constructed 95% CI contained the true estimand value, evaluating the quality of interval estimation. In addition, we estimated Monte Carlo standard errors (MCSEs) using the formulas provided by Morris et al (2019) to quantify the uncertainty in the performance measures due to using a finite number of simulations.morris2019using
@@ 4.6 Results
The complete results for the simulation study are displayed in Tables 1-4. Note that, for the $n=200$ simulations, the MCSEs are less than 0.018 for bias, less than 0.013 for ESE and less than 0.006 for 95% CI coverage; for the $n=1,000$ simulations, the MCSEs are less than 0.011 for bias, less than 0.008 for ESE and less than 0.005 for 95% CI coverage.
@@
Under $n=1000$, all methods perform as expected in terms of bias. The naïve estimator appears biased in all four scenarios (but in KS3 the bias is very small), the DR estimators appear unbiased when either model is correct (Scenarios KS1, KS2 and KS3), and the singly robust estimators avoid bias only when the corresponding propensity score or outcome model is correct. Unlike the three DR methods, the weighted G-computation estimators show bias in Scenario KS2 when the Cauchit link function is used in the outcome model (but not when the canonical logistic link is used). This suggests that the weighted G-computation estimators may be doubly robust for the ATC when the canonical logit link is used, but not otherwise. The trends observed for the bias are similar under $n=200$ with the caveat that the (augmented and non-augmented) weighting estimators that appeared unbiased under $n=1000$, exhibit some small-sample bias in the corresponding scenarios. This is particularly notable in Scenario KS2 and, to a lesser extent, in Scenario KS1, and is probably due to small effective sample sizes after weighting.
@@
Under correct specification of the outcome model (KS1 and KS3), G-computation is the most precise covariate-adjusted estimator, but the augmented estimators are almost as precise (e.g., compare the G-computation estimator which obtains ESE = 0.150 to the DR augmented MAIC estimator which obtains ESE = 0.170 for KS1 with $n=1000$; see Table 1). Moreover, the augmented estimators generally produce precision gains versus their respective non-augmented weighting counterparts. When both the outcome model and the propensity score model are correctly specified (KS1), all augmented estimators have increased precision compared to the non-augmented weighting estimators based on modeling weights, but not necessarily against MAIC (any increase in precision for $n=1000$ is modest). When only the propensity score model is correctly specified (KS2), outcome model misspecification does not induce any meaningful loss in precision for the augmented estimators compared to their non-augmented weighting counterparts.
@@
There have been some concerns in the literature about doubly robust augmented estimators amplifying bias and variance when misspecified weights are combined with a misspecified outcome model.kang2007demystifying Such amplification is not apparent in our simulation study. In KS4, our proposed DR augmented MAIC estimator is the least biased of all estimators and the most precise of the augmented and non-augmented weighting estimators. Under $n=1000$, this advantage is modest when the logit link is used for the outcome model (bias of 0.482 versus 0.512 for G-computation), but more pronounced when the Cauchit link is used (bias of 0.350 versus 0.524 for G-computation).
@@
$n=200$, $\textrm{ATC}=1.116$
1. The naive estimator 0.623 0.326 0.539 1.295
2. IOW 0.027 0.516 0.944 1.979
3. Normalized IOW 0.055 0.456 0.945 1.761
4. MAIC 0.037 0.413 0.962 2.276
With logit link used for outcome model:
5. G-computation 0.019 0.351 0.956 1.431
6. DR augmented IOW 0.029 0.424 0.955 1.663
7. DR augmented normalized IOW 0.029 0.403 0.949 1.605
8. DR augmented MAIC 0.031 0.406 0.955 1.718
9. Weighted G-computation (normalized IOW weights) 0.029 0.401 0.942 1.582
10. Weighted G-computation (MAIC weights) 0.026 0.402 0.946 1.770
With Cauchit link used for outcome model:
5. G-computation 0.006 0.383 0.965 1.689
6. DR augmented IOW 0.026 0.432 0.954 1.682
7. DR augmented normalized IOW 0.026 0.413 0.951 1.635
8. DR augmented MAIC 0.028 0.408 0.957 1.717
9. Weighted G-computation (normalized IOW weights) 0.063 0.914 0.964 2.696
10. Weighted G-computation (MAIC weights) 0.052 0.415 0.963 1.857
$n=1000$, $\textrm{ATC}=1.116$
1. The naive estimator 0.604 0.142 0.008 0.562
2. IOW 0.008 0.205 0.949 0.810
3. Normalized IOW 0.012 0.197 0.942 0.750
4. MAIC 0.008 0.172 0.944 0.660
With logit link used for outcome model:
5. G-computation 0.005 0.150 0.950 0.593
6. DR augmented IOW 0.007 0.175 0.948 0.679
7. DR augmented normalized IOW 0.006 0.173 0.947 0.669
8. DR augmented MAIC 0.007 0.170 0.942 0.653
9. Weighted G-computation (normalized IOW weights) 0.006 0.170 0.940 0.648
10. Weighted G-computation (MAIC weights) 0.007 0.171 0.939 0.649
With Cauchit link used for outcome model:
5. G-computation -0.020 0.158 0.947 0.641
6. DR augmented IOW 0.006 0.178 0.946 0.683
7. DR augmented normalized IOW 0.006 0.175 0.945 0.673
8. DR augmented MAIC 0.006 0.172 0.940 0.658
9. Weighted G-computation (normalized IOW weights) -0.003 0.180 0.948 0.781
10. Weighted G-computation (MAIC weights) -0.004 0.173 0.942 0.677
Table 1: Results from Scenario KS1, where both the logit-link outcome model and the propensity score model are correctly specified.
@@
$n=200$, $\textrm{ATC}=1.215$
1. The naive estimator 0.230 0.324 0.909 1.279
2. IOW 0.023 0.642 0.929 2.393
3. Normalized IOW 0.060 0.511 0.940 1.952
4. MAIC 0.061 0.517 0.965 2.778
With logit link used for outcome model:
5. G-computation 0.088 0.434 0.952 1.739
6. DR augmented IOW 0.050 0.532 0.949 2.093
7. DR augmented normalized IOW 0.053 0.526 0.942 2.007
8. DR augmented MAIC 0.048 0.489 0.953 2.042
9. Weighted G-computation (normalized IOW weights) 0.056 0.480 0.937 1.868
10. Weighted G-computation (MAIC weights) 0.036 0.480 0.949 2.241
With Cauchit link used for outcome model:
5. G-computation 0.222 0.506 0.951 2.085
6. DR augmented IOW 0.046 0.517 0.955 2.053
7. DR augmented normalized IOW 0.046 0.505 0.950 1.974
8. DR augmented MAIC 0.039 0.474 0.958 1.990
9. Weighted G-computation (normalized IOW weights) 0.225 1.950 0.964 5.768
10. Weighted G-computation (MAIC weights) 0.132 0.486 0.972 2.478
$n=1000$, $\textrm{ATC}=1.215$
1. The naive estimator 0.226 0.141 0.658 0.557
2. IOW 0.020 0.283 0.950 1.085
3. Normalized IOW 0.016 0.220 0.939 0.834
4. MAIC 0.015 0.205 0.936 0.779
With logit link used for outcome model:
5. G-computation 0.075 0.188 0.937 0.736
6. DR augmented IOW 0.016 0.228 0.942 0.866
7. DR augmented normalized IOW 0.016 0.224 0.938 0.849
8. DR augmented MAIC 0.014 0.205 0.935 0.776
9. Weighted G-computation (normalized IOW weights) 0.017 0.203 0.936 0.778
10. Weighted G-computation (MAIC weights) 0.014 0.203 0.936 0.770
With Cauchit link used for outcome model:
5. G-computation 0.228 0.284 0.817 1.012
6. DR augmented IOW 0.016 0.224 0.942 0.854
7. DR augmented normalized IOW 0.016 0.220 0.940 0.836
8. DR augmented MAIC 0.012 0.199 0.937 0.756
9. Weighted G-computation (normalized IOW weights) 0.163 1.270 0.935 2.820
10. Weighted G-computation (MAIC weights) 0.114 0.189 0.905 0.743
@@
$n=200$, $\textrm{ATC}=1.068$
1. The naive estimator -0.032 0.307 0.950 1.211
2. IOW 0.147 0.583 0.958 2.246
3. Normalized IOW -0.021 0.397 0.950 1.539
4. MAIC 0.132 0.353 0.955 1.561
With logit link used for outcome model:
5. G-computation 0.015 0.286 0.962 1.180
6. DR augmented IOW 0.028 0.344 0.964 1.404
7. DR augmented normalized IOW 0.027 0.335 0.958 1.338
8. DR augmented MAIC 0.024 0.313 0.962 1.296
9. Weighted G-computation (normalized IOW weights) 0.020 0.305 0.958 1.245
10. Weighted G-computation (MAIC weights) 0.018 0.305 0.959 1.300
With Cauchit link used for outcome model:
11. G-computation -0.031 0.296 0.969 1.316
12. DR augmented IOW 0.008 0.351 0.964 1.412
13. DR augmented normalized IOW 0.005 0.332 0.960 1.348
14. DR augmented MAIC 0.010 0.316 0.965 1.323
15. Weighted G-computation (normalized IOW weights) 0.018 0.701 0.972 2.176
16. Weighted G-computation (MAIC weights) -0.001 0.320 0.970 1.448
$n=1000$, $\textrm{ATC}=1.068$
1. The naive estimator -0.039 0.135 0.937 0.530
2. IOW 0.118 0.231 0.969 0.917
3. Normalized IOW -0.044 0.168 0.933 0.645
4. MAIC 0.104 0.148 0.894 0.574
With logit link used for outcome model:
5. G-computation 0.005 0.124 0.949 0.489
6. DR augmented IOW 0.007 0.144 0.945 0.558
7. DR augmented normalized IOW 0.007 0.142 0.943 0.547
8. DR augmented MAIC 0.006 0.133 0.943 0.517
9. Weighted G-computation (normalized IOW weights) 0.006 0.130 0.943 0.508
10. Weighted G-computation (MAIC weights) 0.006 0.130 0.944 0.506
With Cauchit link used for outcome model:
11. G-computation -0.053 0.124 0.926 0.496
12. DR augmented IOW -0.015 0.144 0.942 0.557
13. DR augmented normalized IOW -0.017 0.141 0.939 0.546
14. DR augmented MAIC -0.010 0.134 0.941 0.520
15. Weighted G-computation (normalized IOW weights) -0.044 0.131 0.937 0.606
16. Weighted G-computation (MAIC weights) -0.045 0.130 0.933 0.518
Table 3: Results from Scenario KS3, where the propensity score model is incorrectly specified. The logit-link outcome model is correctly specified; however, when the Cauchit link is used for the outcome model, both models are incorrectly specified.
@@
$n=200$, $\textrm{ATC}=1.181$
1. The naive estimator 0.521 0.334 0.685 1.330
2. IOW 0.807 0.778 0.912 2.793
3. Normalized IOW 0.582 0.470 0.760 1.840
4. MAIC 0.613 0.463 0.790 2.032
With logit link used for outcome model:
5. G-computation 0.536 0.376 0.750 1.551
6. DR augmented IOW 0.586 0.464 0.767 1.846
7. DR augmented normalized IOW 0.580 0.446 0.753 1.770
8. DR augmented MAIC 0.517 0.407 0.779 1.671
9. Weighted G-computation (normalized IOW weights) 0.546 0.415 0.760 1.692
10. Weighted G-computation (MAIC weights) 0.545 0.423 0.786 1.808
With Cauchit link used for outcome model:
5. G-computation 0.555 0.363 0.799 1.677
6. DR augmented IOW 0.455 0.430 0.850 1.767
7. DR augmented normalized IOW 0.456 0.415 0.837 1.704
8. DR augmented MAIC 0.400 0.386 0.878 1.654
9. Weighted G-computation (normalized IOW weights) 0.575 1.291 0.871 4.635
10. Weighted G-computation (MAIC weights) 0.515 0.405 0.882 2.135
$n=1000$, $\textrm{ATC}=1.181$
1. The naive estimator 0.495 0.146 0.069 0.575
2. IOW 0.778 0.350 0.340 1.420
3. Normalized IOW 0.514 0.197 0.256 0.768
4. MAIC 0.546 0.190 0.174 0.739
With logit link used for outcome model:
5. G-computation 0.512 0.161 0.110 0.636
6. DR augmented IOW 0.536 0.185 0.183 0.729
7. DR augmented normalized IOW 0.534 0.183 0.176 0.716
8. DR augmented MAIC 0.482 0.170 0.196 0.667
9. Weighted G-computation (normalized IOW weights) 0.524 0.179 0.159 0.714
10. Weighted G-computation (MAIC weights) 0.533 0.181 0.162 0.708
With Cauchit link used for outcome model:
5. G-computation 0.524 0.144 0.050 0.578
6. DR augmented IOW 0.395 0.173 0.368 0.679
7. DR augmented normalized IOW 0.400 0.171 0.347 0.666
8. DR augmented MAIC 0.350 0.158 0.403 0.620
9. Weighted G-computation (normalized IOW weights) 0.488 0.165 0.226 1.089
10. Weighted G-computation (MAIC weights) 0.479 0.158 0.155 0.646
@@
Assuming unbiasedness, interval estimation is appropriate if the coverage is approximately equal to 0.95; poor coverage can arise due to bias or to inadequate variance/interval estimation. Coverage is generally close to 0.95 for all covariate adjustment methods in the simulation scenarios under which they are unbiased, which suggests that our proposed non-parametric bootstrap approach for variance estimation is adequate. In the cases in which the DR methods are unbiased (i.e., KS1, KS2 and KS3), coverage rates are between 0.936 and 0.947 across $n=1000$ scenarios. Interestingly, coverage rates seem to increase for these methods under $n=200$ despite the small-sample bias, lying between 0.942 and 0.964. Note that, due to computational limitations, the non-parametric bootstrap approach in the simulation study was conducted with only $B=200$ resamples. This may have impacted the observed coverage rates and we suspect that coverage might be more appropriate when using a larger number of resamples.
While some covariate adjustment methods display bias-induced undercoverage in the scenarios under which they are biased (e.g., MAIC in KS3 under $n=1000$ or all estimators in KS4), they may also display adequate coverage (e.g., normalized IOW with $n=200$) because of excessively large standard errors, probably due to low effective sample sizes after weighting. As observed for KS4, bias-induced undercoverage tends to worsen with higher sample sizes, as interval estimates around the wrong target value become narrower. The naïve estimator displays discernible undercoverage in KS1, KS2 and KS4 (particularly under $n=1000$), not only due to bias but also due to overprecise standard errors that do not account for covariate differences.